Operation Summary for September 5, 2011 - September 7, 2011 |
Monday September 5 The day shift began with Operations monitoring the Tevatron (TEV) store 9061 (132.4E30) and the Recycler stash (197.9E10), with the Antiproton Source (Pbar) stacking (12.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, at 9:18 AM, a Booster RF station (BRF7) tripped off for the first of four times in about three hours. At 10:08 AM, the I- Source vacuum valve closed for six minutes while its vacuum improved. At 10:55 AM, a Linac RF station (LRF3) tripped off for the first of three times in 54 minutes; operators adjusted the station’s gradient set point. During the evening shift, Operations terminated store 9061 at 4:45 PM and put the TEV into HEP shot setup. At 5:46 PM, the TEV quenched at sector DA on a kicker prefire; operators called in kicker personnel. Operators prepared D0 for a two-hour access at 6:18 PM. Kicker personnel reported at 8:30 PM that they had completed their work. Operators put the TEV through a dry squeeze at 8:48 PM. Operators put the TEV through a wet squeeze at 9:38 PM. Operations put the TEV into shot setup at 9:51 PM and established store 9064 at 10:55 PM with an average initial luminosity (AIL) of 203.41E30. Operators reported that LRF1 tripped off a number of times during the shift; it reset. Tuesday September 6 The midnight shift began with Operations monitoring store 9064 (153.2E30) and the stash (72.9E10), with Pbar stacking (9.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:25 AM, BRF7 tripped off for the first of three times in about three hours; it reset. Operators reported at 2:45 AM that the Pelletron's critical device controller had tripped off; operators got everything back to normal. Operators reported problems with the MI-30 BPMs . At 3:36 AM, a MI RF station (MIRF3) tripped off, which caused the NuMI beam permit to trip off; everything reset. Operators tuned on the I- Source and the 750 KeV beamline. MIRF3 tripped off for the first of two times in 30 minutes at 4:23 AM. The NuMI beam permit tripped off for the first of two times in 15 minutes due to an RF indication, but no RF station had tripped off; operators investigated. During the day shift, operators reported at 8:10 AM that Neutron Therapy Facility (NTF) personnel began tuning up for treating one patient later in the morning. At 8:16 AM, the MI-30 BPMs stopped polling; operators contacted a Controls person. Operators reported a possible MI RF LCW leak at MI-60; operators informed RF personnel. Operations terminated store 9064 sometime around 11 AM. Operators prepared CDF and D0 for access at 11:23 AM; they were both out of access by 1:11 PM. Operators put the TEV through a dry squeeze at 1:55 PM and through a wet squeeze at 2:39 PM. Operations put the TEV into shot setup at 3:03 PM. Operators reported at 3:25 PM that the MI-30 BPMs stopped reporting. During the evening shift, Operations established store 9067 at 4:22 PM with an AIL of 389.44E30. At 5:23 PM, BRF5 tripped off and would not reset; a Booster person investigated. Operators bypassed the station and brought on BRF19 in its place. A Controls person reported at 6:36 PM that the MI-30 BPMs were back online. Wednesday September 7 The midnight shift began with Operations monitoring store 9067 (124.02E30) and the stash (210.1E10), with Pbar stacking (19.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:37 AM, the MI beam permit tripped off due to bad beam positions; it reset. At 1:47 AM, MIRF3 and the NuMI beam permit tripped off for the first of six times in two hours and twenty minutes; everything reset. Operators reported at 5:27 AM that the modulator voltage for LRF1 was alarming frequently; operators adjusted its filament current. At 7:07 AM, the Linac beam permit tripped off due to a node polling problem; operators rebooted the node. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac to go down for a power amplifier change on LRF1 after the next store goes in. She will also allow Booster studies, Pbar target blower investigations, and a search for the MI LCW leak; none of these activities required tunnel accesses. |
|
More Information Comments and Suggestions |
| last modified 09/07/2011 email Fermilab |
| Security, Privacy, Legal | |